OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x)= 2x^3 + x^2 – 11x The relative maximum is at (–1.53, 8.3) and the relative minimum is at (1.2, –12.01). The relative maximum is at (–1.53, 12.01) and the relative minimum is at (1.2, –8.3). The relative maximum is at (–1.2, 8.3) and the relative minimum is at (1.53, –12.01). The relative maximum is at (–1.2, 12.01) and the relative minimum is at (1.53, –8

OpenStudy (mathmale):

first of all, will you be using an algebra approach or a calculus approach to answer this question?

OpenStudy (anonymous):

algebra

OpenStudy (mathmale):

In that case your best bet is to graph the function. It'll take a while. But once you've graphed it, you'll be able to spot both the min and the max easily.

OpenStudy (mathmale):

calculus would be much quicker, but you and I may have to save that for later.

OpenStudy (anonymous):

ohhhhh wait

OpenStudy (anonymous):

Its D, right? I did graph it but i didnt know if there was an equation i had to do

OpenStudy (mathmale):

Need to see your graph before I'll offer an opinion on your choice of D. Could you possibly do a screen shot and share it? f(x)= 2x^3 + x^2 – 11x

OpenStudy (anonymous):

Yes, hold on! :)

OpenStudy (mathmale):

Note that f(x)= 2x^3 + x^2 – 11x could be at least partially factored.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute the first derivative of such function

OpenStudy (anonymous):

how/

OpenStudy (michele_laino):

it is simple, we can apply this formula: \[\left( {{x^n}} \right)' = n{x^{n - 1}}\] where at the left side I have indicated the first derivative of \(x^n\) then we can compute the first derivative of the function, term by term. So, we get: \[\left( {2{x^3} + {x^2} - 11x} \right)' = 6{x^2} + 2x\]

OpenStudy (michele_laino):

oops.. i have made an error: \[\Large \left( {2{x^3} + {x^2} - 11x} \right)' = 6{x^2} + 2x - 11\]

OpenStudy (michele_laino):

now, please solve this inequality: \[6{x^2} + 2x - 11 > 0\]

OpenStudy (anonymous):

27 > 0?

OpenStudy (michele_laino):

we have to solve this equation first: \[6{x^2} + 2x - 11 = 0\]

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

27>0 is the answer

OpenStudy (michele_laino):

hint: the solution of such equation are: \[\Large {x_1} = \frac{{ - 1 + \sqrt {67} }}{2},\quad {x_2} = \frac{{ - 1 - \sqrt {67} }}{2}\] so, the solution of the above inequality is: \(x<x_1\) and \(x>x_2\)

OpenStudy (michele_laino):

please note that where \(f′(x)>0\) then function \(f(x)\) is an increasing function

OpenStudy (michele_laino):

so, we have this drawing: 

OpenStudy (mathmale):

If you're attempting to solve this problem graphically, then the graph yuo've shared is perfect for the job. You can see both the min and the max of the given function on this graph. Find the approx coordinates of the min and those of the max. Dani: You say your approach is "algebraic." In this case, graphing is the best approach to finding the min. and the max. of the given function. Michele_Laino's approach uses calculus, which is a powerful tool, but not if you haven't already studied some calculus. any questions?

OpenStudy (michele_laino):

sorry, the solution of the above inequality is: \(x<x_2\) union \(x>x_1\)

OpenStudy (mathmale):

Dani: Once again: Have you studied any calculus? If so, are you familiar with the concept of "derivative?"

OpenStudy (michele_laino):

so I can update the drawing above:  and now, we know what is the point of maximum, and what is the point of minimum